A sample of 100 ages, drawn from a normally distributed population, has a mean of 36 years. The population standard deviation is 4 years. Test the hypothesis that the population mean is 35 years, at = 0.01
1
Expert's answer
2021-07-19T05:47:05-0400

"n=100 \\

\bar{x}=36 \\

\sigma= 4 \\

H_0: \mu=35 \\

H_1: \mu ≠35"

Test-statistic:

"Z = \frac{\bar{x} -\mu}{\sigma/ \sqrt{n}} \\

Z= \frac{36-35}{4 / \sqrt{100}} = 2.5 \\

α=0.01"

Two-tailed test.

Reject H0 if Z≤ -2.575 or Z≥2.575

Conclusion: Since Z= 2.5 < 2.575, accept H0.

There is enough evidence to conclude that the population mean is 35 years, at α= 0.01.